# Scalar-Induced Gravitational Waves in Modified Gravity

### v1

@@ 3.1 $f(R)$ gravity
where $F(R)\equiv\partial f/\partial R$. Eq.  (3.4) can also be recast in terms of the standard Einstein tensor, $G_{\mu\nu}$, by replacing
@@ 3.2 Modified first-order field equations
where $\tau_{\ast}$ and $a_{\ast}$ refer to some reference time.
@@ 3.3 Modified second-order field equations
where $T_{\delta\Phi}(k\tau)$ comes from solving Eq. (3.2). From Eq. (3.2), we note that we cannot use the analytical solution Eq. (2.21) for the first-order scalar perturbation in the GR case, since in pure radiation we would have $R^{(0)}=R^{(1)}=0$. Therefore, we should solve the E.o.M. for $\delta\phi^{(1)}$ using the full GR solution including the sub-dominant matter content in RD. However, for simplicity, we do not include the subdominant matter perturbations when computing the GR part (see Ref. [72] for details on how to include matter perturbations for SIGWs in GR).
@@ 4 Power spectrum of induced GWs
$\langle h_{\lambda_{1}}(\mathbf{k_{1}},\tau)h_{\lambda_{2}}(\mathbf{k_{2}},\tau)\rangle$ $=\frac{1}{a^{2}(\tau)}\int_{\tau_{i}}^{\tau}d\tilde{\tau_{1}}\int_{\tau_{i}}^{\tau}d\tilde{\tau_{2}}G_{\mathbf{k_{1}}}(\tau,\tilde{\tau_{1}})G_{\mathbf{k_{2}}}(\tau,\tilde{\tau_{2}})a(\tilde{\tau_{1}})a(\tilde{\tau_{2}})\times$
@@ 5 Induced GW spectrum in $f(R)$ gravity
where $\rho_{c}(\tau)$ is the critical energy density. Given the expression Eq. (4) for the SIGW power spectrum, we can split the spectral energy density as $\Omega_{\rm GW}(k,\tau)=\Omega_{\rm GW,\text{GR}}(k,\tau)+2\alpha\,\delta\Omega_{\rm GW}(k,\tau)$, where $\Omega_{\text{GR,GW}}$ and $\delta\Omega_{\text{GW}}$ are the GR and beyond-GR contribution, respectively.
@@ 5.1 Numerical calculation of SIGW spectrum in $f(R)$ gravity
From Figures 5.1, 5.2 and 5.3, we observe that the low-frequency tail of the SIGW spectrum gets suppressed by the MG contribution, which thus acts as a damping term. Notice that, for stability reasons, $\alpha$ needs to be positive since it is related to the mass squared of the scalar degree of freedom introduced in $f(R)$ gravity [70]. We argue that observational data from PTA, LISA and LVK could constraint the value of $\alpha$ on scales on which we have limited information. Moreover, even if we are taking $\alpha$ to be large at the time of our GW emission, our model can still agree with current constraints on MG. The discussion about the constraint of $\alpha$ goes beyond the scope of this paper and requires further investigation.

### v2

@@ 3.1 $f(R)$ gravity
where $F(R)\equiv\partial f/\partial R$. The trace of (3.4) gives
$3\Box F(R)+F(R)R-2f(R)=\kappa^{2}T$ (3.5)
where $T=g^{\mu\nu}T_{\mu\nu}$. We know that $f(R)=R$ and $F(R)=1$ in the case of GR without a cosmological constant. Therefore, it is clear from the above equation that $R$ is given directly by $T$ in GR as $R=-\kappa^{2}T$. However, for a general form of $f(R)$, $\Box F(R)\not=0$, thus Eq. (3.5) determines the dynamics of a propagating scalar field $\varphi\equiv F(R)$, the so-called scalaron [86].
Eq.  (3.4) can also be recast in terms of the standard Einstein tensor, $G_{\mu\nu}$, by replacing
@@ 3.2 Modified first-order field equations
where $\tau_{\ast}$ and $a_{\ast}$ refer to some reference time. Eq. (3.2) is the evolution equation for $\delta\phi^{(1)}$, which we will need to solve in order to obtain the source term of the E.o.M. of SIGW. Before proceeding, let us stress an important point. Notice that it is not necessary to use the first-order evolution of the scalaron degree of freedom  [^3] Notice that Eq. (3.5) is nothing but a wave equation for $F(R)$. The field equation (3.5) for the additional propagating scalar degree of freedom at linear order gives $F^{(1)^{\prime\prime}}+2\mathcal{H}F^{(1)^{\prime}}-(\nabla^{2}-a^{2}m^{2}_{s})F^{(1)}=\dfrac{a^{2}\kappa^{2}}{3}T+(4\mathcal{H}\phi^{(1)}\bar{F}^{\prime}+2\phi^{(1)}\bar{F}^{\prime\prime}+\phi^{(1)^{\prime}}\bar{F}^{\prime}+3\psi^{(1)^{\prime}}\bar{F}^{\prime})$ (3.22) where the scalaron mass squared reads $m^{2}_{s}\equiv\dfrac{1}{3}\left(\dfrac{\bar{F}}{\bar{F}_{,R}}-R\right)$, where $\bar{F}_{,R}\equiv d\bar{F}/dR$ (see, e.g., [70, 43]). Notice that in the above equation, no particular form of $f(R)$ has been specified, and no expansion around a spatially flat FLRW space has been performed. to close the system of equations for the fluctuations. This is similar to what happens in Ref. [89]. Nevertheless, we have reported the evolution equation at linear order in footnote (3) for completeness.
@@ 3.3 Modified second-order field equations
where $T_{\delta\Phi}(k\tau)$ comes from solving Eq. (3.2). From Eq. (3.2), we note that we cannot use the analytical solution Eq. (2.21) for the first-order scalar perturbation in the GR case, since in pure radiation we would have $R^{(0)}=R^{(1)}=0$. Therefore, we should solve the E.o.M. for $\delta\phi^{(1)}$ using the full GR solution including the sub-dominant matter content in RD. However, we do not include subdominant matter perturbations in the computation of the GR contribution. This simplification is made to facilitate further calculations, as the cold dark matter correction to the scalar potential in GR contributes only subdominantly to the source term and has a negligible effect on the kernel and resulting energy spectrum. To quantify the impact of this approximation, we compared the numerical kernel including cold dark matter corrections to the one computed without them. The mean percentage difference was found to be as low as $9\times 10^{-10}\%$, with a maximum deviation of only $3.2\times 10^{-7}\%$ across the full range of $x\equiv k\tau$. These negligible differences confirm the validity of neglecting cold dark matter perturbations in the GR kernel within our analysis. A detailed treatment of matter perturbations during the matter-dominated era, including their inclusion in GR, can be found in Ref. [72, 17].
@@ 4 Power spectrum of induced GWs
[…]
@@ 5 Induced GW spectrum in $f(R)$ gravity
where $\rho_{c}(\tau)$ is the critical energy density. Given the expression Eq. (4) for the SIGW power spectrum, we can split the spectral energy density as $\Omega_{\rm GW}(k,\tau)=\Omega_{\rm GW,\text{GR}}(k,\tau)+2\alpha\,\delta\Omega_{\rm GW}(k,\tau)$, where $\Omega_{\text{GR,GW}}$ and $\delta\Omega_{\text{GW}}$ are the GR and beyond-GR contribution, respectively. It is worth noting that, in modified gravity theories such as $f(R)$, the effective Planck mass $M_{\rm eff}^{2}(t)$ can in general be time-dependent (see, e.g., [94]). In the present analysis, the parameter $\alpha$ is treated as a constant. Incorporating a time-dependent $\alpha$, and thus a fully dynamical $M_{\rm eff}^{2}(t)$, is a possible extension of this work and may provide further insights into the evolution of the gravitational wave energy density in $f(R)$ gravity frameworks.
@@ 5.1 Numerical calculation of SIGW spectrum in $f(R)$ gravity
From Figures 5.1, 5.2 and 5.3, we observe that the low-frequency tail of the SIGW spectrum gets suppressed by the MG contribution, which thus acts as a damping term. Notice that, for stability reasons, $\alpha$ needs to be positive since it is related to the mass squared of the scalar degree of freedom introduced in $f(R)$ gravity [70]. We argue that observational data from PTA, LISA and LVK could constraint the value of $\alpha$ on scales on which we have limited information. Moreover, even if we are taking $\alpha$ to be large at the time of our GW emission, our model can still agree with current constraints on MG. The discussion about the constraint of $\alpha$ goes beyond the scope of this paper and requires further investigation. Furthermore, notice that the GR solution is again recovered in the deep infrared regime  [^4] This behavior could also be seen in Fig. 5.3 if the frequency range were extended. and this can be shown analytically as demonstrated in Ref. [42]. Specifically, for $x\ll 1$ (deep infrared) and assuming a primordial spectrum with a sharp peak, the $f(R)$ gravity contribution is suppressed by $\mathcal{O}(\alpha x^{2})$, ensuring that the GR result is always recovered.
To assess which component of the source term in Eq. (3.3) plays a dominant role—the gravitational slip (MG-GR coupling, from the first four lines) or the modified source terms (GR-GR coupling, from the latter lines)—we analyze their respective impacts on the resulting SIGW signal. As shown in Fig. 5.4, the MG-GR coupling consistently exceeds the GR-GR contribution across most of the PTA frequency range. This is confirmed by a pointwise analysis [^5] A pointwise analysis compares the contributions of different terms at each evaluation point rather than using integrated quantities., which shows that the MG-GR term dominates in 62.8% of the sampled points. It’s relative contribution is a median of 87.24% to the total source, indicating that the MG-GR coupling is generally the leading contributor to the modified gravity SIGW signal.
Figure 5.4: Energy density spectrum of gravitational waves in the PTA frequency range, showing the separate contributions from the MG-GR coupling (first four lines in Eq.(3.3)), GR-GR coupling (latter lines of Eq.(3.3)), and the total modified gravity (MG) source. Both axes are plotted on logarithmic scales to highlight features across several orders of magnitude in frequency and amplitude.